OpenStudy (anonymous):

mimi drops a tomato from her window ,which is 100 feet off the ground .the height (that is ,the distance above the ground level ) of the tomato after t seconds is d(t)=100-4t^2. how long does it take the tomato to hit the ground A.10-sq root 5 seconds b.2 seconds c.0 seconds d.5 seconds

OpenStudy (anonymous):

@helper99

OpenStudy (anonymous):

sorry i suck at these type of questions

OpenStudy (anonymous):

do u know anyone that can help

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

do oyu think you an help him?

OpenStudy (anonymous):

its easy

OpenStudy (anonymous):

I don't know this stuff.

OpenStudy (anonymous):

you just plug in the answers into the equation and see which one makes perfect sense

OpenStudy (anonymous):

what is it @FriedRice

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[d(t)=100-4(2)^2 ; d(2) = 84\] so that's not right \[d(t)=100-4(0)^2;d(0) =100\] so that's not right \[d(t)=100-4(5)^2; d(5) = 0\] so that's not right either so answer A must be the right answer

OpenStudy (anonymous):

thats what i believe

OpenStudy (anonymous):

can u help me with some more

OpenStudy (anonymous):

i might be able to but first make a different chat room

OpenStudy (anonymous):

okk